A) On January 1, 2017, Chamberlain Corporation pays $586,000 for a 60 percent ownership in Neville. Annual excess fair-value amortization of $18,500 results from the acquisition. On December 31, 2018, Neville reports revenues of $514,000 and expenses of $359,000 and Chamberlain reports revenues of $775,000 and expenses of $421,000. The parent figures contain no income from the subsidiary. What is consolidated net income attributable to Chamberlain Corporation?

B) Mittelstaedt Inc., buys 60 percent of the outstanding stock of Sherry, Inc. Sherry owns a piece of land that cost $290,000 but had a fair value of $645,000 at the acquisition date. What value should be attributed to this land in a consolidated balance sheet at the date of takeover?
Answer #1

Answer page: 0 chamberlain cosportation pays = $ 586000 for percent -60% fair value amortization = $ 18,500 Srevenues = $5140page 02 = $ 1289000-4798500 = $.490500 less; non controlling interest = (514000 - 359000 - 18500)x 400% 136500 x 40°%. = 1365page:03 the assets and liabilities acquired at their full frin value as on date of acquisition. Hence, in this question milte

Answer page-01 as chamberlain comportation pays = $ 586000 for percent = 60% fair value amortization = $ 18,500 revenues = $ 514000 Euromille Ineville Expenses = ¢ 359000 stevenues = $7450007 Expenses - $ 421000 Schambeylain. (a) consolidated net income attributable to chama - beaglain corporation. combined revenues = $ 514000 + $ 775000 = $ 1889000 combined Expenses = $ 359000 + $ 421000 --$ 480000 fxcess acquisition - date fain value amortization = - 18500 Corobined stevenues - consolidated net Anwme combined Expensest Excess acquisition) =$1289000-($ 480000 +$18500)

page 02 = $ 1289000-4798500 = $.490500 less; non controlling interest = (514000 - 359000 - 18500)x 400% 136500 x 40°%. = 136500 * 0:4 = 54600 2-54600 consolidated net income attributable to the chamber - alocs corporation = consolidated net income. - non-controlling interest 490500 - 54600 = $ 435900 (6) As per this question we need to find out the total value of land to be recognized by miltelstaedt Inc. in its consolidated financial statemente Cxplanation t o for every business combination where 100% Owneg ship is not hold by a single company the acquiren jewgnizes and measwies all identifiable

page:03 the assets and liabilities acquired at their full frin value as on date of acquisition. Hence, in this question miltelstaedt ine. will recognize the land at fair value of $ 645000 The answer is $ 645000